### v1

Abstract
In recent years, consumer Internet of Things (IoT) devices have become widely used in daily life. With the popularity of devices, related security and privacy risks arise at the same time as they collect user-related data and transmit it to various service providers. Although China accounts for a larger share of the consumer IoT industry, current analyses on consumer IoT device traffic primarily focus on regions such as Europe, the United States, and Australia. Research on China, however, is currently rather rare. This study constructs the first large-scale dataset about consumer IoT device traffic in China. Specifically, we propose a fine-grained traffic collection guidance covering the entire lifecycle of consumer IoT devices, gathering traffic from 70 devices spanning 36 brands and 8 device categories. Based on this dataset, we analyze traffic destinations and encryption practices across different device types during the entire lifecycle and compare the findings with the results of other regions. Compared to other regions, our results show that consumer IoT devices in China rely more on domestic services and overally perform better in terms of encryption practices. However, there are still 20/35 devices improperly conduct certificate validation, and 5/70 devices use insecure encryption protocols. To facilitate future research, we open-source our traffic collection guidance and make our dataset publicly available.
keywords:
Chinese Consumer IoT, Traffic Dataset, Traffic Measurement, Privacy, Security
1 Introduction
The consumer Internet of Things (IoT) industry has rapidly developed in recent years, with various devices becoming widespread and offering numerous conveniences. However, the rapid spread of devices has also raised growing concerns regarding security and privacy risks. The data collected by these devices is often transmitted to manufacturers of devices and other service providers. At the same time, data transmission is facing strict legal regulations [1] in China. Therefore, it is necessary to measure consumer IoT traffic from security and privacy perspectives.
Researchers have conducted studies on consumer IoT device traffic in various regions, including Europe [3, 4], the United States [5, 6, 7], and Australia [8, 9], and have identified several security and privacy risks associated with these devices. However, according to IDC [2], China’s consumer IoT industry is projected to reach \$300 billion by 2027. Despite this rapid growth, research on consumer IoT device traffic in China remains limited and lacks comprehensive analysis. This research gap drives us to build a large-scale and comprehensive dataset of consumer IoT traffic in China and conduct a measurement of the dataset.
While constructing the dataset, we introduce methodological improvements compared to existing studies. Previous studies have primarily focused on the device interaction phase [5, 13]. However, we find differences in traffic across different operational phases of the device and limiting the focus on the interaction phase may lead to an incomplete analysis. Therefore, we propose a fine-grained traffic collection guidance that covers the entire lifecycle of the device to address this limitation. The entire lifecycle refers to all phases during normal device usage, including device setup, user interaction, idle, and deletion [10]. Compared with prior works, the traffic we collected shows greater coverage and can support more fine-grained analysis. Furthermore, we open-source our traffic collection guidance and make our dataset publicly available to facilitate further research.
From March 1 to November 29, 2024, we constructed a dataset of consumer IoT traffic in China based on our traffic collection guidance. Furthermore, for devices that can update firmware during the experiment, we also collected traffic after their firmware update to study the differences in device traffic before and after the firmware update. Based on the traffic collected, we analyzed the traffic destinations and encryption practices and then compared our findings with those of the United States and the United Kingdom [5, 13].
Our study reveals that, regarding traffic destinations, Chinese devices show a higher reliance on domestic services. The distribution of traffic destinations for Chinese devices is mainly directed to China (99.8%) and relies on domestic organizations. In contrast, devices from the United Kingdom have a significantly higher proportion of overseas traffic, with a greater reliance on organizations from the United States. Regarding encryption practices, devices from different regions all show good encryption practices, although some issues remain. In particular, Chinese devices have less PII exposure in unencrypted traffic and a lower percentage of unencrypted traffic compared to devices in the United States and the United Kingdom. However, 20/35 devices still use improper certificate verification, and 5/70 use insecure encryption protocols.
Our main contributions include:
  • The first Chinese consumer IoT devices traffic dataset. We constructed a dataset that include 97.5 hours of traffic from 70 devices across 8 categories and 36 brands. We make this dataset publicly available [^1] https://github.com/NKUHack4FGroup/Lifecycle-Based-Traffic-Dataset to facilitate future research.
  • A lifecycle-based fine-grained traffic collection guidance. We propose a fine-grained traffic collection guidance covering the entire lifecycle of the device in a controlled environment. We open-sourced the guidance and are inviting researchers to build a standard, comprehensive, and state-of-the-art consumer IoT dataset together.
  • Security and privacy measurement of Chinese consumer IoT device traffic. We analyze traffic destinations and encryption practices, then compare the results with those from other regions, revealing the characteristics of Chinese consumer IoT devices and regional differences.
2 Background
Consumer IoT devices typically refer to IoT devices designed for personal or household use, such as cameras, speakers, and lights. These devices can be bound to and controlled by the companion apps, following similar operational phases (i.e., lifecycle). During these phases, users can control the device through various methods. Meanwhile, consumer IoT devices typically communicate frequently with cloud servers during operation to support various device functions. This section introduces the concepts and definitions relevant to the lifecycle, control methods, and cloud services mentioned above.
The lifecycle of consumer IoT devices. The lifecycle of consumer IoT devices contains four phases based on a prior work [10]: setup, interaction, idle, and deletion. When users first activate a consumer IoT device, the device will enter the setup phase. During this phase, users need to complete operations such as network configuration and bind the device to a companion app. The interaction phase refers to the phase in which users interact with the device, such as controlling light switches and checking camera status. After setting up the device, we refer to the phase in which the user no longer interacts with the device as the idle phase. Finally, The deletion phase refers to the phase when users unbind the device from the companion app, restoring it to its factory settings.
Control methods of consumer IoT device. Common control methods of consumer IoT devices include four types: Wide Area Network (WAN) control, Local Area Network (LAN) control, physical control, and multimodal control. WAN control refers to scenarios where the controlling smartphone with the companion app and the consumer IoT device are not on the same local network. In this case, commands between the app and the device must be relayed through a cloud server. For example, when a user is away from home during travel, he/she can monitor indoor environments in real-time via a camera through a companion app. LAN control occurs when the controlling smartphone and the consumer IoT device are on the same local network. In this case, communications between the device and the app can be established through local protocols, such as Bluetooth. For example, when a user is at home, he/she can adjust the brightness and color temperature of smart lights through a companion app. Physical control involves direct interaction with the device through physical buttons. For example, users can press physical switch buttons on the device to reset or power it off. Consumer IoT devices with various sensors support multimodal control through different methods. Examples include users waking up speakers through voice commands or interacting with devices via triggers from motion sensors.
Cloud servers of consumer IoT devices. When a consumer IoT device operates, the data it collects is shared with various cloud servers. Following the classification approach used in the study by Ren et al. [5], we categorize these servers into three groups: first party, support party, and third party. First party refers to the manufacturer of the consumer IoT device or the developer of its companion app. Support party refers to organizations that provide computational resources or network services for consumer IoT devices, such as cloud computing, cloud storage service providers, and service providers (e.g., CDN). Third party refers to companies contacted by device and not categorized as first or support parties. We adopt the methodology from Ren et al. [5] to determine server affiliations through a three-step verification process: (1) We first identify the organization of the IP addresses contacted by the device. If the IP’s organization matches the manufacturer or companion app of the consumer IoT device, we classify it as a first party; (2) For unresolved cases, examine official Privacy Policies and Third-Party Data Sharing documents (prioritizing device-specific policies) to identify support party; (3) Categorize any undocumented or unaffiliated servers as the third party.
3 Destination Analysis
For consumer IoT devices in China, traffic directed to overseas locations or third-party destinations is of particular concern, as it may increase security and privacy risks compared to traffic flowing to first-party or support-party servers. This section starts with an analysis of traffic destinations, providing a detailed assessment of security and privacy risks through two key aspects: (1) the geographical distribution of traffic destinations by country and (2) the affiliation analysis of contacted server parties.
3.1 Distribution of Traffic Destinations
This subsection analyzes the country-level distribution of device traffic destinations, focusing on the proportion of traffic to different countries across device types and lifecycle phases.
We first collected IP addresses interacting with devices and queried their geographical affiliations using the IP geolocation API  [12]. Then, we calculate the proportion of traffic going to different countries. In the calculation, we observed that the traffic packet size of cameras, doorbells, and speakers is significantly larger than that of sensors and other devices. If only the proportion of bytes flowing to devices in different countries is counted, there will be a more serious bias. We processed a special calculation method to eliminate the above bias when calculating the traffic proportion.
Let $X$ represent the set of type $A$ device indices in the dataset. For each device $i$ ($i\in X$), let $M_{i}$ denote the traffic volume (in bytes) transmitted to the country $T$, and let $N_{i}$ denote the total traffic volume transmitted to all destinations. The total normalized traffic from type $A$ devices to country $T$ is then given by $A_{T}=\sum_{i\in X}\frac{M_{i}}{N_{i}}$. Further, let $Sum_{A}$ represent the total traffic volume of all type $A$ devices. The proportion of traffic from type $A$ devices directed to country $T$ is computed as: $\frac{A_{T}}{Sum_{A}}$. Similarly, let $Sum_{T}$ represent the total traffic volume from all devices to the country $T$, and $Sum_{all}$ denote the total traffic volume from all devices to all destinations. The overall proportion of traffic directed to the country $T$ across all devices is then given by: $\frac{Sum_{T}}{Sum_{all}}$.
We used Chiplot [34] to generate a Sankey diagram (Figure 1) to visualize the distribution of traffic destinations. In terms of destination, 99.8% of all device traffic remained within China, with only 0.2% directed overseas. From the perspective of device types, hubs, humidifiers, lights, sensors, and plugs did not show overseas traffic, likely due to their simple communication with cloud servers. A total of 27 devices showed overseas traffic, all cameras, speakers, or doorbells. For each type of device, the cameras had the highest proportion of foreign traffic at 0.29%, followed by speakers at 0.05% and doorbells at 0.07%.
Figure 1: Proportion of device traffic routed to different countries, grouped by device type and traffic destination country.
The United States is the most common overseas destination, accounting for 0.11% of total traffic. In particular, dns.google.com is the most frequently accessed foreign domain, communicated by 20 devices. Other domains involved in United States traffic include ntp.com, amazonaws.com, and iotcplatform.com, supporting services such as time synchronization (e.g., 1.cn.pool.ntp.org), remote access based on TUTK (e.g., all-d-master-tutk.iotcplatform.com, used by Xiaobai doorbells and Jooan cameras), and cloud computing (e.g., p2ps2.coolkit.cn, accessed by Sonoff cameras). Traffic to other countries accounted for less than 0.1% of the total and involved 18 domain names. Among them, 10 domains (e.g., ntp.org, time.cloudflare.com) are related to NTP services. Other services included Google Cloud (e.g., bc.googleusercontent.com), IoT platform management (e.g., all-master.iotcplatform.com), IP address lookup (e.g., api.ipify.org), and AWS cloud hosting (e.g., eu-central-1.compute.amazonaws.com). It should be noted that 27 devices contacted NTP servers in 12 countries and 7 devices contacted NTP servers in multiple countries. This increases the security risk of being affected by international cyber attacks. In contrast, domestic NTP services could offer safer and more stable connections.
Figure 2: Overseas traffic destinations of devices grouped by lifecycle phases. Device Lifecycle Phases: SU (setup), ID (idle), IR (interaction), DE (deletion). The destinations with values less than 0.1 are magnified.
To analyze the distribution of traffic across the device lifecycle phases, we visualized the proportional traffic destination during the different lifecycle phases in Figure 2. Since over 99% of the traffic terminates within China, the domestic traffic is omitted from the figure. The results show that during the idle phase, cameras, speakers and doorbells exhibit a relatively high proportion of overseas traffic, and these traffic primarily flow to NTP services. During the setup phase, the cameras and speakers show a notable amount of overseas traffic and involve diverse services, such as DNS service, IP address lookup, IoT platform management, and NTP service. This means the device configurations rely on more types of overseas traffic. During both the interaction and deletion phases, the frequency of overseas destinations decreases. This likely suggests that the main functionalities of IoT devices are built on domestic services.
Additionally, for 10 devices that updated firmware, none of them generated overseas traffic before or after the firmware updates.
3.2 Service Contacted by Devices
During device operation, frequent traffic communication occurs between devices and cloud service providers. We calculated the number of cloud server affiliations contacted by devices in different lifecycle phases. Table  1 shows the average number of service providers contacted by devices of each type.
Table 1: The Average Number of Servers Contacted by Devices in Different Lifecycle phases.
Life Cycle Server Camera Doorbell Hub Humidifier Light Plug Sensor Speaker
First Party 3.14 2.33 0.56 2.00 1.33 0.83 1.33 3.29
Setup Support Party 1.96 1.00 0.78 0.00 0.33 0.67 0.00 1.14
Third Party 1.18 0.67 0.00 0.00 0.00 0.00 0.00 0.86
First Party 1.82 1.67 0.44 1.00 1.33 1.08 2.33 2.29
Idle Support Party 2.54 1.00 0.67 0.00 0.33 0.50 0.00 1.57
Third Party 1.29 1.33 0.00 0.00 0.00 0.00 0.33 1.57
First Party 1.43 4.00 0.33 1.00 0.67 0.58 1.33 2.14
Interaction Support Party 1.71 1.33 0.22 0.00 0.17 0.25 0.00 1.71
Third Party 0.93 1.67 0.00 0.00 0.00 0.00 0.33 1.29
First Party 0.43 2.33 0.33 1.00 0.67 0.58 1.00 0.57
Deletion Support Party 0.57 0.00 0.33 0.00 0.17 0.17 0.00 0.29
Third Party 0.21 1.00 0.00 0.00 0.00 0.00 0.33 0.29
First Party 4.18 4.33 1.11 2.00 1.83 1.75 3.67 5.29
Total Support Party 4.89 2.33 1.00 0.00 0.50 0.83 0.00 3.86
Third Party 3.93 2.00 0.00 0.00 0.00 0.00 0.33 3.00
First, regardless of the type of device, first party contacts constituted a significant proportion of all device communications. Second, the interaction with the support party is closely related to the device’s computational resource needs. Among the eight device types, the camera, doorbell, and speaker interact more frequently with the support party. As a result, many manufacturers choose to collaborate with cloud service providers to ensure the devices can leverage cloud computing and storage capabilities, thereby enhancing their stability. The support parties contacted by these devices typically provide cloud computing and storage services (e.g., oss-cn-beijing.aliyuncs.com, ap-beijing.myqcloud.com), video streaming services (e.g., video.qq.com, stream.qqmusic.qq.com), and network communication services (e.g., dns.alidns.com).
Compared with first and support parties, third party domains primarily involve data sharing, advertising services, geolocation, and personalized recommendation services. For example, the Huawei speaker contacts statist.tingmall.com for music streaming services, which may be related to third-party audio components used by the device; the IMOU camera contacts api.weathercn.com for weather query services. Additionally, we also observe that many devices access seemingly unrelated third-party domains. For instance, Xiaomi speaker, Konka camera, Lenovo camera, Pisen camera, and Jooao camera all access the domain www.baidu.com. These observations suggest that while third party services offer various personalized features, they are not explicitly mentioned in privacy policies. As a result, unexpected destination contacts may raise user privacy concerns.
Additionally, for 10 devices support firmware updates, there was no significant change in the number of third party service providers contacted. Overall, traffic destinations showed little change after the firmware updates, indicating that device firmware updates have little impact on traffic destinations in the short term.
Commonly Contacted Organizations. To further assess whether the devices’ communications exhibit excessive reliance on specific organizations, we analyzed the organization of traffic destinations and ranked them based on interaction frequency. We focused on the companies most frequently contacted by consumer IoT devices and listed the top ten companies in Table 2. According to the table, 20 devices in the dataset contacted Google to use its DNS service. In contrast, the local DNS provider Greatbit, is used by only 17 devices, which is fewer than Google.
In real-time communication for consumer IoT devices, stable and efficient DNS services are essential. While Google Public DNS is commonly used due to its global network and high availability, GreatBit, as a domestic provider, offers lower latency and better stability for local devices. However, the analysis shows that Google DNS is used more frequently than expected. In consumer IoT device real-time communication scenarios, stable and efficient DNS services are crucial. As a domestic service provider, GreatBit offers lower latency and higher stability for local devices, making it more suitable to meet the demands of domestic devices. Nevertheless, the analysis shows that compared with the local DNS service, the frequency of the device using Google DNS is higher. In addition to Google and Greatbit, the Organizations most frequently contacted by devices include Alibaba Cloud, Xiaomi, and Tencent. The services provided by these companies through their domains include cloud storage services (e.g., oss-cn-shenzhen.aliyuncs.com), NTP services (e.g., time1.cloud.tencent.com, ntp.aliyun.com), and video streaming and media content (e.g., findermp.video.qq.com, vv.video.qq.com), inditcating that devices rely on these companies for the computational resources they provide.
Table 2: Organizations Commonly Contacted by Devices.
Organization Google Alibaba Cloud Xiaomi Greatbit Tencent Baidu Huawei Tuya NST Jooan
Frequency 20 18 17 17 11 10 9 5 4 3
3.3 Summary
By analyzing traffic destinations, we find that most Chinese devices rely on domestic services. Specifically, 99.8% of the traffic is directed to domestic servers that depend on domestic organizations. This means that Chinese IoT devices feature a high degree of localization and greater autonomy in control. From the perspective of lifecycle phases, we can infer that the main functionalities of IoT devices, e.g., controlling and monitoring, are built on demostic services. Most foreign services used by Chinese consumer IoT devices are basic Internet services like NTP and DNS, which can also be replaced by Chinese service providers. In addition to contacting first parties, many devices from different manufacturers commonly communicate with major domestic organizations such as Xiaomi, Alibaba, and Tencent. This allows them to receive information from many devices and potentially obtain the capability to construct user profiles, e.g., inferring types of device in a household.
4 Encryption Analysis
To ensure the security of consumer IoT device traffic, devices often encrypt the communication. However, encryption implementations vary across brands and device types, with differing security levels. In this section, we measure encryption practices through 4 key aspects: (1)the proportion of encrypted traffic, (2) the encryption protocols, (3) the certificate security in TLS, and (4) the exposure of personally identifiable information (PII) in unencrypted traffic.
4.1 Encrypted Traffic Proportion Analysis
Consumer IoT device communications may involve sensitive information such as user personal data and surveillance audio or video. Transmitting such data in plaintext raises interception, tampering, and unauthorized risks. Thus, we leverage the entropy of the traffic to identify and analyze the proportion of unencrypted communications.
Common traffic analysis tools, such as Wireshark, can easily distinguish between encrypted protocols like TLS and QUIC, but cannot clearly classify customized encryption protocols. Therefore, we adapted the method proposed by Ren et al. [5] to determine whether traffic is encrypted by calculating the payload entropy of the packets. First, for HTTP protocol packets, we classify them into text, compressed, or media based on the context_type field. For SSL protocol, if the payload entropy is greater than 0.8, it is classified as encrypted. For DNS protocol, packets containing the dns_key field are treated as encrypted. For packets that cannot be classified in the first step, we further classify them based on the magic_number of the payload. The magic_number refers to a unique identifier or pattern of bytes in a packet that helps to determine its type or format, often used to identify specific file types or data structures. If the packet is identified as compressed or media data, it is categorized as compressed or media. Finally, for the remaining packets, we perform a final classification based on entropy values: packets with entropy greater than 0.9 are classified as encrypted, those with entropy less than 0.4 are classified as text, and those with entropy between 0.4 and 0.9 are classified as unknown. Ultimately, we classify the traffic into encrypted traffic, unencrypted traffic, and unknown traffic.
Figure 3 presents a heat map visualization illustrating the proportional distribution of encrypted, unknown, and unencrypted traffic across various lifecycle phases of consumer IoT devices.
Figure 3: For each device type, the heatmaps depict the average percentages of encrypted, unknown, and unencrypted traffic transmitted across respective lifecycle phases.
The heatmaps show that the proportion of unencrypted traffic is generally low, with no device type exceeding 8.5% of unencrypted traffic. This indicates that encrypted communication dominates the data transmission of most devices. However, significant variation exists across device types. For instance, camera, doorbell, and speaker devices exhibit an average encrypted traffic share of over 60%, largely driven by the security and privacy requirements in audio and video data transmission. In contrast, the humidifier and plug devices have lower average encrypted traffic proportions of 36% and 38%, while their unknown traffic proportions are the highest among all devices, at 50% and 49%, respectively. Analysis of traffic entropy reveals that most of these devices’ traffic falls within the entropy range of 0.4-0.7. Upon further inspection, it appears that these devices do not use standard encryption protocols but instead rely on proprietary protocols developed by the manufacturers. These proprietary protocols have unstable entropy. By reverse engineering, an network attacker could readily deduce the syntax and semantics of the manufacturer’s protocols, thereby enabling them to manipulate and monitor users’ data. When considering the lifecycle phases of the device, the proportion of encrypted traffic is relatively higher during the setup and interaction phases than in the other two phases. These phases involve crucial services such as device initialization and configuration, cloud computing, and media data transmission. Therefore, more encrypted traffic in these phases ensures communication security.
Additionally, for 10 devices that support firmware updates, Most devices exhibited increased encrypted traffic proportions after updates. For example, TP-Link Camera increased the average encrypted traffic proportion from 65.75% pre-update to 72.56% post-update. This indicates that the encryption communication is becoming increasingly prevalent.
4.2 Encryption Protocol Analysis
In addition to standard encryption protocols such as TLS and SSL, some manufacturers use proprietary encryption protocols based on UDP [13]. Therefore, we first check whether standard encryption protocols, such as SSL or TLS, are used for encrypted traffic. Encrypted traffic not using these protocols is classified as encrypted with proprietary protocols. In the analysis, we statistic the number of devices using each encryption protocol with the results presented in Table 3.
Table 3: Distrubution of Encryption Protocols Deployed in Devices.
Protocol TLS 1.2 TLS 1.1 TLS 1.3 SSL SSLv2 Other Protocols
Device Num 58 1 3 5 3 10
Through analysis, we found that TLS 1.2 is widely used for encrypted communication across devices, with 58 devices using TLS 1.2 at least once. However, some devices (e.g., IMOU doorbells, Pisen cameras, and Tenda cameras) continue to rely on insecure SSL or TLS 1.1 protocols, which accounts for 7.6% of the total communication traffic on average. Analyzing the destination IP addresses of these traffic showed that most targets are telecommunications service providers (e.g., China Mobile) or cloud service providers (e.g., Alibaba Cloud). Moreover, not all encrypted communications from these devices use SSL or TLS 1.1. There is a mixed use with other protocols (e.g., TLS 1.2). This may result from manufacturers neglecting to fully deprecate outdated protocols during certain functionality upgrades, leaving some connections still using SSL or TLS 1.1. On the other hand, while TLS 1.3 offers significant improvements in performance and security, only three devices in our dataset use it. This indicates that despite its clear advantages, its adoption remains limited, and manufacturers must make greater efforts to implement the latest security protocols. Additionally, we identified 10 devices (e.g., Aqara Camera, Aqara Plug, Xiaovv Camera) using proprietary protocols developed by the manufacturers. The security of these protocols remains to be further analyzed.
In conclusion, although some devices still use outdated encryption protocols and custom encryption formats present security risks, TLS 1.2 remains the most widely used encryption protocol, and the adoption of TLS 1.3 requires more attention from manufacturers in future updates.
Additionally, for the 10 devices that support firmware updates, most used TLS 1.2 for encryption before the update, while three devices relied on proprietary protocols developed by the manufacturers. The XiaobaiY2 camera also used SSL in addition to TLS 1.2. After the update, all devices upgraded their encryption protocols. In addition to continuing to use TLS 1.2, they also adopted TLS 1.3. The three devices that previously used proprietary protocols switched to TLS 1.2, and the XiaobaiY2 camera stopped using SSL, now relying only on TLS 1.2 and TLS 1.3 for encrypted communication.
4.3 TLS Certificate Security Analysis
The TLS 1.2 protocol is the most widely used encryption protocol in our dataset. To further understand its security, this section presents a statistical analysis of the certificate for the 58 devices in the dataset that use the TLS 1.2 protocol. The results show that 43 devices use a public key certificate model for communication. The remaining 15 devices used the Pre-Shared Key (PSK) model without relying on certificates. Given the crucial role digital certificates play in TLS encryption security, we further analyzed the self-signed certificate usage and the encryption algorithms’ security. Additionally, we conducted dynamic tests using the mitmproxy to assess the validity of certificates when devices powered on again.
4.3.1 Certificate Usage
Overall, the certificate security of domestic devices is generally good, but some insecure certificates were still identified. Table 4 statistics the number of TLS implementations that may be insecure, including insecure signature algorithms, insecure cryptographic algorithms, and self-signed certificates. Among the 43 devices, 7 use insecure algorithms, such as the SHA1 signature algorithm and RSA-1024 cryptographic algorithm, primarily for communication with Aliyun and AWS servers. These algorithms are now considered insecure and vulnerable to attacks. Secondly, approximately 39% of the devices use self-signed certificates in at least one TLS 1.2 connection. These certificates are typically signed by CA authorities like ClickPKI or self-signed by the device manufacturers. Many self-signed certificates have excessively long validity periods, with one Lenovo camera certificate valid until 2120. According to current security standards, the validity period of TLS certificates should not exceed 13 months [14]. As pointed out by Dong et al. [15], the lack of public oversight for long-term vendor-signed certificates poses potential vulnerabilities, as they may not meet the stringent standards of trusted public CAs, thereby increasing the security risks within the consumer IoT ecosystem.
Table 4: TLS 1.2 implementations of different device type.
TLS 1.2 implementation Camera Doorbell Plug Hub Light Sensor Speaker
Insecure Signature Algorithms 3 0 1 1 0 0 2
Insecure Cryptographic Algorithms 1 0 0 1 0 0 0
Self-signed Certificates 13 1 0 1 1 1 0
4.3.2 Certificate Validation
Ideally, devices should check whether the certificates from servers are valid. If a forged certificate is detected, the device should either return an error message or reject the TLS connection. However, if a device ignores certificate anomalies and proceeds to establish a connection, the certificate validation mechanism may have vulnerabilities.
To investigate this issue, we first successfully captured traffic during device reconnections and discovered that 35 devices would conduct TLS connection with certificate exchange with 92 servers during the reconnection period. Next, we conducted a mitmproxy certificate replacement attack on these TLS connections. The results are presented in Table 5.
Table 5: Error messages or device behavior after the certificate replacement
Error Message and Device Behavior Device Number Server Number
“Unknown CA” 11 17
“Decrypt error” 2 2
“Bad certificate” 2 2
“Close notify” 5 15
“Decode error” 1 1
Disconnects and reconnects 12 19
Server handshake failed 1 2
Unable to connect to the internet 1 1
Communicates normally 20 33
The results show that approximately 64% of the servers responded with errors, such as ”Unknown CA,” ”Decrypt error,” ”Close notify” alerts, or actively terminated the connection. Meanwhile, 33% of the servers associated with 20 devices did not verify the certificates. By decrypting the traffic of these servers, we were able to obtain sensitive information such as device logs, configuration details, and authentication credentials. This indicates that many consumer IoT devices follow insecure certificate validation practices, leaving them vulnerable to certificate replacement attacks.
4.4 PII Exposure in Unencrypted Traffic
Personally Identifiable Information (PII) refers to any data that can identify an individual [16]. Common PII includes names, ID numbers, phone numbers, email addresses, etc. Unencrypted traffic containing PII significantly elevates the risk of user data exposure. In this section, we identify PII in unencrypted traffic using text matching and regular expressions to measure the security practice. To comprehensively analyze PII exposure, we input as much personal information as possible into the companion app when we collect traffic.
The analysis reveals that the exposure of PII in unencrypted communication traffic is minimal in traffic. Only a small amount of PII (including the names of five well-known singers) is detected in the traffic from four speaker devices, and no inputted personal information is found. These singer names appeared in the song requests during the interaction phase. Although this information does not directly reveal the user’s identity, it could still be exploited by malicious third parties to infer user preferences or deliver targeted ads, thereby posing a risk of privacy leakage and data misuse.
4.5 Summary
This section measures the encryption practices of consumer IoT devices. The study reveals that Chinese manufacturers adopt safe encryption protocols for device communications, and encrypted traffic dominates most communications. However, some devices still use deprecated protocols, though their traffic constitutes a small proportion(7.6%), and some devices use proprietary protocols developed by the manufacturers rather than standardized protocols. bringing about certain security risks. Encouragingly, following firmware updates, 10 devices adopted more secure encryption protocols after the update, indicating a positive trend toward encryption protocol usage. Most devices use secure certificates, but prevalent self-signed certificates and insecure algorithms persist as potential risks. In terms of certificate verification, many devices have the problem of insecure certificate verification mechanisms, which may also bring potential risks. While direct PII exposure in unencrypted traffic is rare, inferred user preferences could enable targeted advertising. These findings indicate that the devices should adopt standard encryption protocols and reduce reliance on self-signed certificates to further minimize risks.
5 Comparisons on Different Regions
This section compares consumer IoT device traffic measurement results in China with other regions. We choose Ren et al.[5] and Paracha M. T. et al.[13] for comparison, as their studies align with our research purpose, focusing on traffic destinations and encryption analysis.
5.1 Traffic Destinations
In terms of traffic destinations, compared to consumer IoT devices in the United States and the United Kingdom, Chinese devices predominantly communicate domestically. According to Ren et al. [5], the 35 devices in the United Kingdom communicated with 7 overseas regions. Approximately 70% of the United Kingdom device traffic flowed outside their regions (primarily to the United States). In contrast, although Chinese consumer IoT devices contacted 18 overseas countries, overseas traffic only accounted for 3% of total bytes. This highlights Chinese consumer IoT devices’ stronger reliance on domestic services than their United Kingdom counterparts.
Regarding commonly contacted organizations, Chinese consumer IoT devices predominantly interact with domestic providers. In the Ren et al. study [5], the United Kingdom devices are mainly connected to the United States organizations, such as Amazon, Google, and Akamai. In our analysis, except for Google, Chinese devices primarily communicated with domestic organizations such as Alibaba Cloud, Xiaomi, and Tencent. These providers offer diverse services, including IoT platform services, cloud storage services, and video streaming and media content (See Section 3.2 for details).
At last, according to the findings of Ren et al. [5], the number of third parties contacted by devices is generally low, with the setup phase involving more third parties than other phases, likely due to devices establishing initial connections during this phase. In our study, the number of third parties contacted by devices is also low, suggesting that devices in different regions tend to communicate more with first and support parties. On the other hand, we find that the number of third parties contacted during the idle phase is higher than in other phases, which differs from Ren et al.’s findings. Devices in the United States and the United Kingdom tend to complete third party dependency initialization during the setup phase, while Chinese devices are more likely to maintain long-term connections with third parties during the idle phase for purposes such as data uploading, status synchronization, or push services.
5.2 Encryption Analysis
Encryption practices vary regionally, but unencrypted traffic proportions remain low overall. Ren et al.’s. [5] proposed entropy-based encryption detection, reporting unencrypted traffic proportions below 13% for devices of the United States and the United Kingdom devices. Using similar methods, we come to a similar conclusion: the percentage of unencrypted traffic of consumer IoT devices in China is low, with a peak of 8.5%.
Regarding encryption protocol usage, TLS 1.2 remains the most widely adopted, while TLS 1.3 adoption is still limited. Paracha M. T. et al. [13] analyzed 40 consumer IoT devices in the United States in 2021 and found that TLS 1.2 is the dominant protocol, with minimal use of TLS 1.3 and deprecated versions. Our findings align with these results. Among 70 Chinese devices, 82% utilized TLS 1.2, while 7% still relied on insecure protocols (TLS 1.1/SSL), and 14% employed proprietary protocols developed by manufacturers. For certificate validation, Paracha M. T. et al. [13] examined 32 devices and found that 11 were vulnerable to certificate replacement. In our study, 15 out of 35 devices generated error messages when certificate replaced, while 20 devices exhibited no errors or anomalies, indicating that certificate verification remains a security concern.
In terms of PII exposure, Ren et al.’s study [5] identified PII in unencrypted traffic, including MAC addresses, UUIDs, device IDs, geolocation data, and user-assigned device names. In contrast, Chinese devices exposed minimal PII—primarily singer names in speaker’s traffic, indirectly revealing music preferences. This indicates stronger privacy protection in Chinese devices’ unencrypted traffic.
5.3 Summary
A clear regional difference is observed in the distribution of traffic destinations and encryption practices. Chinese consumer IoT devices are more reliant on local service providers, in contrast to devices in the United Kingdom, which predominantly rely on international cloud platforms such as Amazon. This suggests that China has developed a more self-reliant and locally integrated IoT ecosystem, with limited dependence on international platforms. Encryption practices and certificate implementations are generally consistent across regions, ensuring relative security, but a certain proportion of devices still use outdated protocols across regions, which could pose potential security and privacy risks. Meanwhile, there are weaknesses with certificate verification in devices from different regions, posing a security concern. The PII contained in the unencrypted traffic of Chinese consumer IoT devices is less than that of their counterparts in the United Kingdom and the United States.
6 Related Work
The Measurement of IoT traffic. Researchers have conducted a lot of measurements on IoT traffic, mainly focusing on security and privacy. Security measurements involve mapping IoT backend servers [17], analyzing the impact of botnets on devices [26, 27], and investigating data leakage and device vulnerabilities through companion apps [29, 31, 32, 33]. Said et al. [17] mapped IoT backend servers by analyzing ISP passive traffic data, revealing relationships between backend providers. Arman et al. [26] assessed the impact of two ISP security policies on the Mirai botnet. Stephen et al. [27] studied the Hajime botnet’s effect on devices using active scanning and passive DNS backscatter traffic collection. These studies primarily focus on device-level evaluations while overlooking the user perspective. Privacy measurements include traffic destination identification [30], user data collection analysis [23, 24], and the effectiveness of existing privacy protection policies [25]. Mandalari et al. [30] proposed an automated method to identify IoT traffic destinations, distinguishing between essential and non-essential ones. Aniketh et al. [23] examined privacy leaks in local communication models using standard protocols. Dubois et al. [24] investigated how, when, and why smart speakers record external audio, while Anna Maria et al. [25] tested the effectiveness of existing privacy protection measures. Compared with these studies, our work directly analyzes IoT device security practices by evaluating traffic destinations and encryption practices. Ren et al. [5] and Paracha et al. [13] have similar research purposes with ours. Muhammad Talha Paracha et al. [13] evaluated the security of TLS connections in IoT devices, exploring how these connections evolve over time and whether certificates are properly validated. Ren et al. [5] analysis the information exposure of devices in the United Kingdom and the United States. Inspired by their work, we introduce an analysis of Chinese devices and compare the result with other regions.
Name Area1 Source2 Categories Number(IoT)3 Period Size Time Life cycle4
SU ID IR DE
Ours [6] US SC 10 8 2020.3 11.5GB 11 days
YourThings [7] US SC 15 45 2018.3 233GB 13 days
IoTDNS [19] US SC 28 53 2019.8 366MB 2 months
UNSW [8] AUS SC 17 28 2016.10 9.72GB 6 months
BoT-IoT [9] AUS SL 5 5 2018.4 69.3GB 2 months
Mon(IoT)r [5] US&UK SC 15 81 2018.9 12.9GB -
PingPong [11] US SC 12 19 2019 40.3GB 51 days
HomeSnitch [20] US SC 13 57 2021.3 595MB 8 days
IoT_Sentinel [3] FI SC 6 31 2016 61.4MB -
IoT23 [4] CZ SC 3 3 2018 21GB 1 year
N-BaIoT [21] IL SC 3 9 2018.3 240GB -
NSL-KDD [22] US SL - - 1998.5 4.06MB 7 weeks
CN-CIoT CN SC 8 70 2024.3 2.7GB 3 months
Table 6: The Summary of IoT Device Datasets.
  • 1
    US (United States), UK (United Kingdom), AUS (Australia), FI (Finland), CZ (Czech Republic), IL (Israel), CN (China).
  • 2
    Data Collection Method Labels:SC (Self-Collection), CR (Crowdsourcing), SL (Device Simulation).
  • 3
    The number of IoT devices in the dataset.
  • 4
    Device Lifecycle Phases: SU (setup), ID (idle), IR (interaction), DE (deletion).
IoT Dataset Construction. The IoT community has developed diverse datasets to facilitate researches in consumer IoT devices. Table 6 summarizes the most high-cited datasets. Among self-collected datasets, Mon(IoT)r [5] is the most cited and largest in device count, followed by UNSW  [8], YT  [18], Ours [6], and PingPong  [11]. Previous studies have provided valuable datasets for consumer IoT device research, but there are still some limitations. First, these datasets are collected before 2021, lacking recent consumer IoT traffic. Second, most existing datasets focus on device traffic from Europe and the United States, while traffic from Asia is rare. Third, current datasets do not fully consider the entire lifecycle of devices during the collection, restricting the analysis from the perspective of the lifecycle. In contrast, our dataset includes traffic with the entire lifecycle of the device, enabling researchers to conduct more fine-grained analyses of consumer IoT device traffic.
In summary, we construct the first Chinese consumer IoT traffic dataset that covers the entire lifecycle of a device, completing the research gap in the dataset for China. This dataset enables a more fine-grained analysis of traffic destinations and encryption practices and compares our measurements with those of other regions. By constructing the dataset and conducting measurements, this study provides new perspectives and insights for future research.
7 Conclusion
This study constructs the largest dataset of consumer IoT device traffic in China, proposes a fine-grained traffic collection guidance covering the entire lifecycle of the device, and conducts measurements on traffic destination and encryption practice.
We collected 97.5 hours of traffic based on 70 consumer IoT devices across 8 categories and 36 brands. Through analysis, we draw the following key conclusions. First, 99.8% of traffic from Chinese consumer IoT devices is routed domestically, with only 0.2% directed overseas. In contrast, approximately 70% of the United Kingdom device traffic flowed overseas regions. This indicates minimal reliance on foreign services. However, a notable observation is that most devices depend on services from major corporations like Alibaba Cloud, Xiaomi, and Tencent, indicating that most devices rely on the services of these organizations, and these organizations may also obtain more user information from devices. Regarding encryption practices, Chinese consumer IoT devices demonstrate generally robust performance. The unencrypted traffic of Chinese devices peaked at 8.5%, lower than the other regions, and Chinese devices have less PII exposure in unencrypted traffic. Nevertheless, some devices still employ deprecated encryption protocols, insecure TLS certificate algorithms (e.g., SHA1, RSA-1024), and self-signed certificates. More critically, during device certificate replacement, certain devices exhibited an absence of certificate validation mechanisms, exposing communication to potential security and privacy risks.
We open-source our traffic collection guidance and make our dataset publicly available on website [^2] https://github.com/NKUHack4FGroup/Lifecycle-Based-Traffic-Dataset to facilitate future research.
[…]
This work was partly supported by the National Natural Science Foundation of China No.62102198 and the Key Program of the National Natural Science Foundation of China under No.62032012 and No.62432012.